OpenStudy (anonymous):

A bike travel 7 ft in 1/3 of a second. At the same speed, how many ft will it travel in 5 second

OpenStudy (mikeymaximum):

I think that the answer is 105

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

Ok that's what I got. Just want to make sure I was doing it right. Thanks :)

OpenStudy (texaschic101):

7/(1/3) = x/5 --- 7 ft to 1/3 sec = x ft to 5 sec this is a proportion so we will cross multiply 1/3x = 35 -- divide both sides by 1/3 x = 35/(1/3) x = 35 * 3/1 x = 105 thats what I got